OpenStudy (anonymous):

minimize: 6x-27/5x x 2/4x-18 i'm stuck in this stupid reduction to the lowest terms kinda problem.

OpenStudy (anonymous):

\[\frac{6x-27}{5x}\times \frac{2}{4x-18}\]?

OpenStudy (anonymous):

factor as \[\frac{3(2x-9)\times 2}{5x\times 2(2x-9)}\] then you see some stuff cancel

OpenStudy (anonymous):

am i suppose to reply

OpenStudy (anonymous):

does the (2x -9) cancel out

OpenStudy (anonymous):

yes it does cancel out

OpenStudy (anonymous):

There is something else that cancels out also

OpenStudy (anonymous):

is the answer 3/5x

OpenStudy (anonymous):

yes it is